Manuel is 43 years old. Delaney is 30 years old. How many years ago was Manuel's age 2 times Delaney's age?

Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

Manuel is 43 years old. Delaney is 30 years old.
How many years ago was Manuel's age 2 times Delaney's age?
:
let y = no. of years for this to be true
43 - y = 2(30 - y)
distribute the 2
43 - y = 60 - 2y
-y + 2y = 60 - 43
y = 17 years ago
:
:
See if that is true
43 - 17 = 26
30 - 17 = 13 which is half of 26

Answer by Alan3354(69443)

Manuel is 43 years old. Delaney is 30 years old. How many years ago was Manuel's age 2 times Delaney's age?
----
43 - 30 = 13
--> 13 & 26
etc